OpenStudy (anonymous):

Alan can chop 3/4 of a carrot in 1/8 min. How many carrots can he chop in 1 min?

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@zoejay1

jabez177 (jabez177):

\[\frac{ 3 }{ 4 }\times \frac{ 1 }{ 8 }\]

jabez177 (jabez177):

3 x 1 is?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

@zoejay1

jabez177 (jabez177):

4 x 8 is?

OpenStudy (anonymous):

32

jabez177 (jabez177):

\[\frac{ 3 }{ 32 }\times \frac{ 1 }{ 1 }\] because we're supposed to divide but dividing fractions is the same as multiplying their reciprocal.

jabez177 (jabez177):

Wait... I may be mistake, let me check again.

jabez177 (jabez177):

Can you put a fraction in your choice? If no then I made a mistake.

OpenStudy (anonymous):

would it be 6???

jabez177 (jabez177):

How did you get it?

OpenStudy (anonymous):

http://calculatorsoup.com